OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6? ANSWERS BELOW PLS HELP

OpenStudy (anonymous):

slope = -5; y-intercept = (0, -6) slope = five thirds; y-intercept = (0, 2) slope = -three fifths; y-intercept = (0, 2) slope = five thirds; y-intercept = (0, -6)

OpenStudy (anonymous):

Solve for y. Once you do: \(y=mx+b\) \(m=\)slope\) \(b = y\)-intercept

OpenStudy (anonymous):

The y-intercept is the point where x = 0, so if you plug in 0 for x: \(y=m(0)+b \implies y=b\)

OpenStudy (shamim):

5 x-3 y=-6 3 y=5 x+6 y=(5/3) x+6/3 y=(5/3)x+2 now u hv to compare this equation with y=mx+c so slope m=5/3 and y intercept=2